OpenStudy (anonymous):

Calculate dy/dx Simplify your answer. y=x^9sqr of x

OpenStudy (anonymous):

use the product rule

OpenStudy (anonymous):

must be the second tuesday of the week already...

OpenStudy (anonymous):

man, they havent deleted me yet

OpenStudy (anonymous):

its there fault im still on the juice

OpenStudy (anonymous):

use \[(fg)'-f'g+g'f\] with \[f(x)=x^9, f'(x)=9x^8, g(x) =\sqrt{x}, g'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

i can quit any time. really. i can. it was just the bar was open and...

OpenStudy (anonymous):

thanks guys i got 19x^17/2/2